Choose and exponential function that represents a real-world situation and graph the function. Analyze the graph.
Answer by rothauserc(4718)

The following exponential function is used to calculate the amount of money after a certain time
:
A = P(1 + r/n)^tn, where A is the amount, P is the principle, r is the rate, n is the number of times the interest is compounded in a year, t is the number of years
:
Suppose we want to answer the following question
:
How much money should be invested at 5% compounded quarterly for 20 years so that you have $20000 at the end of the 20 years?
:
20000 = P(1 + 0.05/4)^(20*4)
:
20000 = P(2.701484941)
:
P = 20000 / 2.701484941 = 7403.335735206
:
We need to invest $7403.34
: